How do you solve this three variable equation?
d + e - f = 11
f + d + 5 = e
2e - 12 = f

Answer by ankor@dixie-net.com(22740)

: How do you solve this three variable equation?
d + e - f = 11
f + d + 5 = e
2e - 12 = f
:
Arrange all in the same order
d + e - f = 11
d - e + f = -5
0 +2e - f = 12
:
Multiply the 1st equation by - 1
-d - e + f = -11
d - e + f = -5
0 + 2e - f = 12
--------------------adding eliminates d and e, find f
0 + 0 + f = -4
f = -4
:
Find e using the last equation
2e - f = 12
2e - (-4) = 12
2e + 4 = 12
2e = 12 - 4
2e = 8
e = +4
:
Find d using the 1st equation
d + e - f = 11
d + 4 (-4) = 11
d + 4 + 4 = 11
d = 11 - 8
d = 3
:
Check solutions in the 2nd equation
d - e + f = -5
3 - 4 - 4 = -5